On October 28, 2015, the DJIA opened at 17,716.11. The divisor at that time was 0.149677385. In October 2015, Goldman Sachs was the highest priced stock in the DJIA and Cisco was the lowest. The closing price for Goldman Sachs on October 27, 2015, was $181.99, and the closing price for Cisco was $28.09. Calculate the difference between the new DJIA levels under the following two scenarios (i.e.g, Scenario 1 - Scenario 2). Scenario 1: Suppose the next day the other 29 stock prices remained unchanged and Goldman Sachs increased 5 percent. Scenario 2: Suppose the next day the other 29 stock prices remained unchanged and Cisco Sachs increased 5 percent. (Round your answer to 2 decimal places.)
Answer #1

DJIA on 28th October 2015 = 17,716.11

Divisor at that time = 0.149677385

Let us remove the effect of the divisor,

=17,716.11?*0.149677385?= 2651.70102

Rounding off to 2 decimals = 2651.70 ………. (1)

?Scenario 1:

Goldman Sachs on October 27, 2015 = $181.99

  • 5% Increase in Goldman Sachs with other 29 stocks having no change :

= $181.99 * 5% = 9.9095 ………..(2)

  • Index value without divisor effect after increase (1) + (2)

= 2651.70 + 9.9095

= 2661.6095 Which is 2661.61 (Approx) ……… (3)

  • DJIA in Scenario 1 = (3) * Divisor

= 2661.61/0.149677385

= 17782.31 ……….(4)

Scenario 2:

If Cisco Increased 5% on next day:

Cisco on October 27, 2015 = $28.09

  • 5% Increase in Cisco with other 29 stocks having no change :

= $28.09 * 5% = 1.445 ………..(5)

  • Index value without divisor effect after increase (1) + (5)

= 2651.70 + 1.445

= 2653.145 Which is 2653.15 (Approx) ……… (6)

  • DJIA in Scenario 1 = (6) * Divisor

= 2653.15/0.149677385


= 17725.79 ……….(7)

Difference between the New DJIA levels under both the Scenarios (4)-(7) =

= 17782.31 - 17725.79

= 56.52 points


?